If a temperature increase from 22.0 ∘C to 35.0 ∘C triples the rate constant for a reaction, what is the value of the activation barrier for the reaction?
Answer #1

ln(k2/k1) = - Ea/R (1/T2 - 1/T1)

K2 /K1 = 3

Ea = ?

R = 8.314 JK-1Mole -1

T2 = 35 +273 = 308 K

T1 = 22+273 = 295 K

Substitute the values in the equaiton and find Ea

ln 3 = -Ea /8.314 (1/308-1/295)

1.10 = Ea x 1.72 x 10-5

Ea = 1.1 x 105 /1.72

Ea = 6.395 x 104 Jouls

Ea (Activation barrier) = 6.395 x 104 Joules /mol or 63 kilo Joules / mole